OpenStudy (anonymous):

Solve of x in the proportion 3x/3x-36 = -1/x A. x=3 and x=-4 B. x=-3 and x=4 C. -3 D. -12

OpenStudy (anonymous):

3x/(3x - 36) = -1/x 3x/(3(x - 12)) = -1/x x/(x - 12) = -1/x x^2/(x-12) = -1 x^2 = -(x - 12) x^2 + x - 12 = 0 (x + 4)(x - 3) = 0 thus (x + 4) = 0 or (x - 3) = 0 equivalently, x = -4 or x = 3 A.